# A canonical Hamiltonian formulation of the Navier-Stokes problem

### 14 Dec 2023 v2

@@ Abstract
This paper presents a novel Hamiltonian formulation of the isotropic Navier-Stokes problem based on a minimum-action principle derived from the principle of least squares. This formulation uses the velocities $u_{i}(x_{j},t)$ and pressure $p(x_{j},t)$ as the field quantities to be varied, along with canonically conjugate momenta deduced from the analysis. From these, a conserved Hamiltonian functional $H^{*}$ satisfying Hamilton’s canonical equations is constructed, and the associated Hamilton-Jacobi equation is formulated for both compressible and incompressible flows. This Hamilton-Jacobi equation reduces the problem of finding four separate field quantities ($u_{i}$,$p$) to that of finding a single scalar functional in those fields—Hamilton’s principal functional $\text{S}^{*}[t,u_{i},p]$. Moreover, the transformation theory of Hamilton and Jacobi now provides a prescribed recipe for solving the Navier-Stokes problem: Find $\text{S}^{*}$. If an analytical expression for $\text{S}^{*}$ can be obtained, it will lead via canonical transformation to a new set of fields which are simply equal to their initial values, giving analytical expressions for the original velocity and pressure fields. Failing that, if one can only show that a complete solution to this Hamilton-Jacobi equation does or does not exist, that will also resolve the question of existence of solutions. The method employed here is not specific to the Navier-Stokes problem or even to classical mechanics, and can be applied to any traditionally non-Hamiltonian problem.
@@ 1.1   A motivating example
with unique solution $v(t)=e^{-t}$. Here $v(t)$ can be interpreted as the velocity of a lumped mass moving in a viscous medium in one dimension with linear damping. Like the traditional Navier-Stokes equations [7, 8, 9, 10, 11, 12, 13], this too is an intrinsically non-Hamiltonian problem, in that there is no action $\mathcal{S}$ for which Hamilton’s principle ($\delta\mathcal{S}=0$) yields the governing equation $\dot{v}=-v$. And yet, if we simply differentiate both sides of the equation ($\ddot{v}=-\dot{v}$), use the original equation to write $\dot{v}=-v$, and apply the additional initial condition $\dot{v}(0)=-v(0)=-1$, we arrive at the mathematically equivalent second-order problem
@@ 1.2   The Navier-Stokes problem
where $\rho$ is the constant and uniform density, $u_{i}=u_{i}(x_{j},t)$ is the velocity field, $p=p(x_{j},t)$ is the pressure field, $b_{i}=b_{i}(x_{j},t)$ is the body force field, subscript Roman indices label Euclidean tensor components ($i,j=1,2,3$), the $x_{j}$ are Eulerian spatial coordinates, $t$ is time, $\mu$ is the viscosity, a dot over a symbol denotes a partial time derivative ($\dot{u}_{i}=\partial{u}_{i}/\partial t$), a comma in a subscript indicates a spatial gradient ($p_{,i}=\partial p/\partial x_{i}$), and we employ the Einstein summation convention on repeated subscript indices. In the case of a uniform gravitational field, $b_{i}=g_{i}$ coincides with the local acceleration due to gravity; however, in what follows, we make no assumptions about the functional form of $b_{i}(x_{j},t)$: it is completely arbitrary. There are four unknown field quantities: $u_{i}(x_{j},t)$ and $p(x_{j},t)$.
@@ 2   Literature review
(Here Millikan [1] uses the notation $u=u_{1}$, $v=u_{2}$, and $w=u_{3}$.) More generally, it has been shown that the existence of variational formulations is related to self-adjointness of the system with respect to a standard duality relation, a property that all non-conservative systems lack [67]. Within the last 80 years, many alternative methods have been developed in an attempt to circumvent the non-self-adjointness of dissipative systems [68, 69, 2, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 56, 57, 80, 81]. The mathematical study of alternative variational methods as applied to the Navier-Stokes equations in particular remains an ongoing endeavor [82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 18].
@@
Gonzalez and Taha [93] and Taha and Gonzalez [94] have recently applied Gauss’s principle of least constraint [95] to the Navier-Stokes problem. It should be noted that Gauss’s principle [95] is fundamentally different from Hamilton’s principle [4, 5, 6]. Whereas the Hamiltonian framework involves an invariant action integral and employs variations in the coordinates (or, in continuum mechanics, the field quantities), Gauss’s principle employs variations in the accelerations. As a result, the framework of Gauss’s principle does not lead to canonical transformations.
@@ 3   Lagrangian formulation of the problem
Although we are primarily interested in the incompressible form of the equations given by (16) and (17), for reasons that will become clear later, here we will begin with the compressible form of the equations, with the understanding that we will eventually take the incompressible limit. For the compressible case, the linear momentum balance and continuity equations are given by
$\mathcal{R}_{i}[t,x_{j},u_{j},p,\rho]\equiv\rho\dot{u}_{i}+\rho u_{i,j}u_{j}+p_{,i}-\mu u_{i,jj}-(\mu+\lambda)u_{j,ji}-\rho b_{i}=0,$ (21)
@@ 3.1   Second-order formulation
$\mathcal{R}_{i}[t,x_{j},u_{j},p]\equiv\hat{\rho}\dot{u}_{i}+\hat{\rho}u_{i,j}u_{j}+p_{,i}-\mu u_{i,jj}-(\mu+\lambda)u_{j,ji}-\hat{\rho}b_{i}=0,$ (24)
@@
*$\mathcal{L}^{*}[t,x_{j},u_{j},p]=\frac{1}{2}\mathcal{R}_{i}\mathcal{R}_{i}+\frac{1}{2}\mathcal{R}_{4}\mathcal{R}_{4}.$ (27)
@@ 3.2   Equivalence of the first- and second-order formulations
The proof is straightforward. Consider the two formulations separately, and impose on the second-order formulation identical auxiliary conditions to those of the first-order formulation. In particular, just like the simple example given in Section 1.1, the second-order formulation requires additional auxiliary conditions over and above those applied to the first-order formulation. These include initial conditions making $\mathcal{R}_{i}(x_{j},0)=0$ and $\mathcal{R}_{4}(x_{j},0)=0$ for all $x_{j}\in\mathcal{V}\cup\partial\mathcal{V}$, along with boundary conditions making $\mathcal{R}_{i}(x_{j},t)=0$, $\mathcal{R}_{i,j}(x_{j},t)=0$, and $\mathcal{R}_{4}(x_{j},t)=0$ for all $x_{j}\in\partial\mathcal{V}$ and all times $t$. By supposition, the auxiliary conditions applied to the two formulations are identical, so it suffices to show that $(u_{i},p)$ satisfies the governing field equations (24), (25) of the first-order formulation ($\mathcal{R}_{i}=0$ and $\mathcal{R}_{4}=0$) everywhere in $\mathcal{V}$ and at all times if and only if $(u_{i},p)$ satisfies the Euler-Lagrange equations (29), (30) of the second-order formulation everywhere in $\mathcal{V}$ and at all times.
@@
Conversely, suppose that $(u_{i},p)$ satisfies the Euler-Lagrange equations (29), (30) of the second-order formulation everywhere in $\mathcal{V}$ and at all times. We note that $(\mathcal{R}_{i}=0,\mathcal{R}_{4}=0)$ constitutes an equilibrium solution of the Euler-Lagrange equations (29), (30). Thus, because the initial conditions are chosen such that $\mathcal{R}_{i}(x_{j},0)=0$ and $\mathcal{R}_{4}(x_{j},0)=0$ for all $x_{j}\in\mathcal{V}\cup\partial\mathcal{V}$, and because the boundary conditions are chosen such that $\mathcal{R}_{i}(x_{j},t)=0$, $\mathcal{R}_{i,j}(x_{j},t)=0$, and $\mathcal{R}_{4}(x_{j},t)=0$ for all $x_{j}\in\partial\mathcal{V}$ and all times $t$, then $\mathcal{R}_{i}$ and $\mathcal{R}_{4}$ will remain identically zero everywhere in $\mathcal{V}$ for all future times. Thus, $(u_{i},p)$ satisfies the governing field equations (24), (25) of the first-order formulation everywhere in $\mathcal{V}$ and at all times. This completes the proof, and we have established that the two formulations are equivalent. $\square$
@@ 4   Hamiltonian formulation of the problem
*$\mathcal{H}^{*}[t,x_{i},u_{i},p,\pi_{i},\pi_{4}]=$ $\frac{1}{2}\frac{1}{(\hat{\rho})^{2}}{\pi_{i}\pi_{i}}-\frac{1}{\hat{\rho}}\biggl{(}\hat{\rho}u_{i,j}u_{j}+{p_{,i}-\mu u_{i,jj}-(\mu+\lambda)u_{j,ji}}-\hat{\rho}b_{i}\biggr{)}\pi_{i}$
@@
We return now to our previous observation concerning the vanishing of the residuals. While $H^{*}$ vanishes for the particular fields $(u_{i},p)$ that satisfy the governing field equations (24), (25) of the first-order formulation, it does not vanish for every conceivable $(u_{i},p)$. The latter would imply, according to Equations (42), that $\dot{u}_{i}\equiv 0$ and $\dot{p}\equiv 0$, which is not generally the case. This observation, and the fact that Equations (43) faithfully reproduce the Euler-Lagrange equations (29), (30) of the second-order formulation, confirm that the Hamiltonian formulation described above is, in fact, a legitimate reformulation of the problem. [In the following section, we develop the Hamilton-Jacobi theory as it relates to the present formulation, the goal being to find a canonical transformation to a new set of fields ($\phi_{i}$, $\phi_{4}$) and conjugate momenta for which the Hamiltonian does vanish identically.]
@@
*$\mathcal{H}^{*}[t,x_{i},u_{i},p,\pi_{i}]=\frac{1}{2}\frac{1}{{\rho}^{2}}{\pi_{i}\pi_{i}}-\frac{1}{{\rho}}\biggl{(}{\rho}u_{i,j}u_{j}+p_{,i}-\mu u_{i,jj}-{\rho}b_{i}\biggr{)}\pi_{i},\quad\hat{\rho}^{\prime}\equiv 0,$ (45)
@@ 4.1   Hamilton-Jacobi equation
In the present context, the role of Hamilton’s principal function is played by a characteristic functional $\text{S}^{*}=\text{S}^{*}[t,u_{i},p]$ (not to be confused with the action $\mathcal{S}^{*}$, although they are related; see Appendix A), which is the solution to the following Hamilton-Jacobi equation:
${H}^{*}\left[t,u_{i},p,\frac{\delta\text{S}^{*}}{\delta u_{i}},\frac{\delta\text{S}^{*}}{\delta p}\right]+\frac{\partial\text{S}^{*}}{\partial t}=0,$ (46)
@@
We conclude this section by remarking that, in the inviscid limit ($\mu=\lambda=0$), all of the preceding formalism remains perfectly well-posed. In that limit, the present approach yields a mathematically equivalent second-order formulation of the inviscid Euler equations, as one would expect.
@@ 6   Case Study
Hamilton’s principal functional $\text{S}^{*}=\text{S}^{*}[t,u_{1},p]$ can be expressed as an integral over $x_{2}$ only, since the other spatial coordinates do not appear and may be integrated out. In this way, we may write the Hamilton-Jacobi equation as follows:
@@ 7   Conclusion
Perhaps the most obvious route would be to solve the Hamilton-Jacobi equation—either (47) for the compressible case or (48) for the incompressible case—for Hamilton’s principal functional $\text{S}^{*}[t,u_{i},p]$ directly. If a complete solution for $\text{S}^{*}$ can be found, it will lead via canonical transformation to a new set of fields which are equal to their initial values, thereby giving analytical expressions for the original velocity and pressure fields. Or, failing that, if one can simply establish based on emerging analytical techniques that a complete solution to this Hamilton-Jacobi equation does (or does not) exist under the usual assumptions, that will also settle the question of existence of solutions.

### 02 Feb 2024 v3

@@ Abstract
This paper presents a novel Hamiltonian formulation of the isotropic Navier-Stokes problem based on a minimum-action principle derived from the principle of least squares. This formulation uses the velocities $u_{i}(x_{j},t)$ and pressure $p(x_{j},t)$ as the field quantities to be varied, along with canonically conjugate momenta deduced from the analysis. From these, a conserved Hamiltonian functional $H^{*}$ satisfying Hamilton’s canonical equations is constructed, and the associated Hamilton-Jacobi equation is formulated for both compressible and incompressible flows. This Hamilton-Jacobi equation reduces the problem of finding four separate field quantities ($u_{i}$,$p$) to that of finding a single scalar functional in those fields—Hamilton’s principal functional $\text{S}^{*}[u_{i},p,t]$. Moreover, the transformation theory of Hamilton and Jacobi now provides a prescribed recipe for solving the Navier-Stokes problem: Find $\text{S}^{*}$. If an analytical expression for $\text{S}^{*}$ can be obtained, it will lead via canonical transformation to a new set of fields which are simply equal to their initial values, giving analytical expressions for the original velocity and pressure fields. Failing that, if one can only show that a complete solution to this Hamilton-Jacobi equation does or does not exist, that will also resolve the question of existence of solutions. The method employed here is not specific to the Navier-Stokes problem or even to classical mechanics, and can be applied to any traditionally non-Hamiltonian problem.
@@ 1.1   A motivating example
with unique solution $v(t)=e^{-t}$. Here $v(t)$ can be interpreted as the velocity of a lumped mass moving in a viscous medium in one dimension with linear damping. Like the traditional Navier-Stokes equations [7, 8, 9, 10, 11, 12, 13], this too is an intrinsically non-Hamiltonian problem, in that there is no action $\mathcal{S}$ for which Hamilton’s principle ($\delta\mathcal{S}=0$) yields the governing equation $\dot{v}=-v$. Yet if we simply differentiate both sides of the equation ($\ddot{v}=-\dot{v}$), use the original equation to write $\dot{v}=-v$, and apply the additional initial condition $\dot{v}(0)=-v(0)=-1$, we arrive at the mathematically equivalent second-order problem
@@ 1.2   The Navier-Stokes problem
where $\rho$ is the constant and uniform density, $u_{i}=u_{i}(x_{j},t)$ is the velocity field, $p=p(x_{j},t)$ is the pressure field, $b_{i}=b_{i}(x_{j},t)$ is the body force field, subscript Roman indices label Euclidean tensor components ($i,j=1,2,3$), the $x_{j}$ are Eulerian spatial coordinates, $t$ is time, $\mu$ is the dynamic viscosity, a dot over a symbol denotes a partial time derivative ($\dot{u}_{i}=\partial{u}_{i}/\partial t$), a comma in a subscript indicates a spatial gradient ($p_{,i}=\partial p/\partial x_{i}$), and we employ the Einstein summation convention on repeated subscript indices. To be clear, the notation $u_{i}(x_{j},t)$ signifies that each component of the velocity field is a function of all three spatial coordinates $(x_{1},x_{2},x_{3})$ and time $t$. It has the same meaning as other common notations, such as $u_{i}(\mathbf{x},t)$ and $\mathbf{u}(\mathbf{x},t)$. Likewise for all other field quantities. In the case of a uniform gravitational field, $b_{i}=g_{i}$ coincides with the local acceleration due to gravity; however, in what follows, we make no assumptions about the functional form of $b_{i}(x_{j},t)$: it is completely arbitrary. There are four unknown field quantities: $u_{i}(x_{j},t)$ and $p(x_{j},t)$.
@@ 2   Literature review
(It should be noted that the six terms referred to above come from the convective acceleration $u_{i,j}u_{j}$, and Millikan [1] uses the notation $u=u_{1}$, $v=u_{2}$, and $w=u_{3}$.) More generally, it has been shown that the existence of variational formulations is related to self-adjointness of the system with respect to a standard duality relation, a property that all non-conservative systems lack [67]. Within the last 80 years, many alternative methods have been developed in an attempt to circumvent the non-self-adjointness of dissipative systems [68, 69, 2, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 56, 57, 80, 81]. The mathematical study of alternative variational methods as applied to the Navier-Stokes equations in particular remains an ongoing endeavor [82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 18].
@@
Gonzalez and Taha [93], Taha and Gonzalez [94], and Taha, Gonzalez, and Shorbagy [95] have recently applied Gauss’s principle of least constraint [96] to the Navier-Stokes problem. Using Gauss’s principle, Taha, Gonzalez, and Shorbagy [95] have shown that, for an incompressible fluid, the magnitude of the pressure gradient is minimum over the domain, which they term the Principle of Minimum Pressure Gradient (PMPG). When applied to an inviscid fluid in two dimensions, the PMPG provides a closure condition based in first principles that could be considered a generalization of the Kutta condition to smooth geometries. It should be noted that Gauss’s principle [96] is fundamentally different from Hamilton’s principle [4, 5, 6]. Whereas the Hamiltonian framework involves an invariant action integral and employs variations in the coordinates (or, in continuum mechanics, the field quantities), Gauss’s principle employs variations in the accelerations. As a result, the framework of Gauss’s principle does not lead to canonical transformations.
@@ 3   Lagrangian formulation of the problem
Although we are primarily interested in the incompressible form of the equations given by (16) and (17), here we will begin with the compressible form of the equations, with the understanding that we will eventually take the incompressible limit. For the compressible case, the linear momentum balance and continuity equations are given by
$\mathcal{R}_{i}[u_{j},p,\rho;x_{k},t]\equiv\rho\dot{u}_{i}+\rho u_{i,j}u_{j}+p_{,i}-\mu u_{i,jj}-(\mu+\lambda)u_{j,ji}-\rho b_{i}=0,$ (21)
@@ 3.1   Second-order formulation
$\mathcal{R}_{i}[u_{j},p;x_{k},t]\equiv\hat{\rho}\dot{u}_{i}+\hat{\rho}u_{i,j}u_{j}+p_{,i}-\mu u_{i,jj}-(\mu+\lambda)u_{j,ji}-\hat{\rho}b_{i}=0,$ (24)
@@
*$\mathcal{L}^{*}[u_{j},p;x_{k},t]=\frac{1}{2}\mathcal{R}_{i}\mathcal{R}_{i}+\frac{1}{2}\mathcal{R}_{4}\mathcal{R}_{4}.$ (27)
@@ 3.2   Equivalence of the first- and second-order formulations
The proof is straightforward. Consider the two formulations separately, and impose on the second-order formulation identical auxiliary conditions to those of the first-order formulation. In particular, just like the simple example given in Section 1.1, the second-order formulation requires additional auxiliary conditions over and above those applied to the first-order formulation. These include initial conditions making $\mathcal{R}_{i}(x_{k},0)=0$ and $\mathcal{R}_{4}(x_{k},0)=0$ for all $x_{k}\in\mathcal{V}\cup\partial\mathcal{V}$, along with boundary conditions making $\mathcal{R}_{i}(x_{k},t)=0$, $\mathcal{R}_{i,j}(x_{k},t)=0$, and $\mathcal{R}_{4}(x_{k},t)=0$ for all $x_{k}\in\partial\mathcal{V}$ and all times $t$. By supposition, the auxiliary conditions applied to the two formulations are identical, so it suffices to show that $(u_{i},p)$ satisfies the governing field equations (24), (25) of the first-order formulation ($\mathcal{R}_{i}=0$ and $\mathcal{R}_{4}=0$) everywhere in $\mathcal{V}$ and at all times if and only if $(u_{i},p)$ satisfies the Euler-Lagrange equations (29), (30) of the second-order formulation everywhere in $\mathcal{V}$ and at all times.
@@
Conversely, suppose that $(u_{i},p)$ satisfies the Euler-Lagrange equations (29), (30) of the second-order formulation everywhere in $\mathcal{V}$ and at all times. We note that $(\mathcal{R}_{i}=0,\mathcal{R}_{4}=0)$ constitutes an equilibrium solution of the Euler-Lagrange equations (29), (30). Thus, because the initial conditions are chosen such that $\mathcal{R}_{i}(x_{k},0)=0$ and $\mathcal{R}_{4}(x_{k},0)=0$ for all $x_{k}\in\mathcal{V}\cup\partial\mathcal{V}$, and because the boundary conditions are chosen such that $\mathcal{R}_{i}(x_{k},t)=0$, $\mathcal{R}_{i,j}(x_{k},t)=0$, and $\mathcal{R}_{4}(x_{k},t)=0$ for all $x_{k}\in\partial\mathcal{V}$ and all times $t$, then $\mathcal{R}_{i}$ and $\mathcal{R}_{4}$ will remain identically zero everywhere in $\mathcal{V}$ for all future times. Thus, $(u_{i},p)$ satisfies the governing field equations (24), (25) of the first-order formulation everywhere in $\mathcal{V}$ and at all times. This completes the proof, and we have established that the two formulations are equivalent. $\square$
@@ 4   Hamiltonian formulation of the problem
*$\mathcal{H}^{*}[u_{i},p,\pi_{j},\pi_{4};x_{k},t]=$ $\frac{1}{2}\frac{1}{(\hat{\rho})^{2}}{\pi_{i}\pi_{i}}-\frac{1}{\hat{\rho}}\biggl{(}\hat{\rho}u_{i,j}u_{j}+{p_{,i}-\mu u_{i,jj}-(\mu+\lambda)u_{j,ji}}-\hat{\rho}b_{i}\biggr{)}\pi_{i}$
@@
We return now to our previous observation concerning the vanishing of the residuals. While $H^{*}$ vanishes for the particular fields $(u_{i},p)$ that satisfy the governing field equations (24), (25) of the first-order formulation, it does not vanish for every conceivable $(u_{i},p)$. The latter would imply, according to Equations (42), that $\dot{u}_{i}\equiv 0$ and $\dot{p}\equiv 0$, which is not generally the case. This observation, and the fact that Equations (43) faithfully reproduce the Euler-Lagrange equations (29), (30) of the second-order formulation, confirm that the Hamiltonian formulation described above is, in fact, a legitimate reformulation of the problem. In the following section, we develop the Hamilton-Jacobi theory as it relates to the present formulation, the goal being to find a canonical transformation to a new set of fields ($\phi_{i}$, $\phi_{4}$) and conjugate momenta for which the Hamiltonian does vanish identically.
@@
*$\mathcal{H}^{*}[u_{i},p,\pi_{j};x_{k},t]=\frac{1}{2}\frac{1}{{\rho}^{2}}{\pi_{i}\pi_{i}}-\frac{1}{{\rho}}\biggl{(}{\rho}u_{i,j}u_{j}+p_{,i}-\mu u_{i,jj}-{\rho}b_{i}\biggr{)}\pi_{i},\quad\hat{\rho}^{\prime}\equiv 0,$ (45)
@@ 4.1   Hamilton-Jacobi equation
In the present context, the role of Hamilton’s principal function is played by a characteristic functional $\text{S}^{*}=\text{S}^{*}[u_{i},p,t]$ (not to be confused with the action $\mathcal{S}^{*}$, although they are related; see Appendix A), which is the solution to the following Hamilton-Jacobi equation:
${H}^{*}\left[u_{i},p,\frac{\delta\text{S}^{*}}{\delta u_{j}},\frac{\delta\text{S}^{*}}{\delta p};t\right]+\frac{\partial\text{S}^{*}}{\partial t}=0,$ (46)
@@
We conclude this section by remarking that, in the inviscid limit ($\mu=\lambda=0$), all of the preceding formalism remains perfectly well-posed. In that limit, the present approach yields a mathematically equivalent second-order formulation of the inviscid Euler equations, as one would expect. Interested readers will find the full details in Appendix B.
@@ 6   Case study
Hamilton’s principal functional $\text{S}^{*}=\text{S}^{*}[u_{1},p,t]$ can be expressed as an integral over $x_{2}$ only, since the other spatial coordinates do not appear and may be integrated out. In this way, we may write the Hamilton-Jacobi equation as follows:
@@ 7   Conclusion
Perhaps the most obvious route would be to solve the Hamilton-Jacobi equation—either (47) for the compressible case or (48) for the incompressible case—for Hamilton’s principal functional $\text{S}^{*}[u_{i},p,t]$ directly. If a complete solution for $\text{S}^{*}$ can be found, it will lead via canonical transformation to a new set of fields which are equal to their initial values, thereby giving analytical expressions for the original velocity and pressure fields. Alternatively, if one can simply establish based on emerging analytical techniques that a complete solution to this Hamilton-Jacobi equation does (or does not) exist under the usual assumptions, that will also settle the question of existence of solutions.
@@
By no means is either of the above programs trivial. As pointed out in Section 4.1, solving equations containing Volterra [23] functional derivatives is itself a long-standing open problem in mathematics, which has received very little attention since the first half of the twentieth century [107, 108, 109, 110, 111, 112, 113]. One might even go so far as to call it a “forgotten” open problem (as did one of the reviewers of the present paper, who generously drew our attention to References [108, 109, 110, 111, 112, 113]). We see the lack of work on such equations as a challenge, yes, but at the same time we also see it as a significant opportunity for advancing the field of analytical continuum mechanics. Perhaps, despite an apparent increase in complexity, a rigorous theory of such equations can be developed after all, in which case the present formulation of the Navier-Stokes problem might be solved as one example. We submit that, at the very least, such an endeavor merits further study, which we intend to continue in future work.